X+5(x-1)=19
2x=10(x-4)
6x-11=-2x+5
18x-15=-21+12x
1/2(10x-6)=x-15
7(1-x)=-4x-11
3(2x+12)=-2(x-4)+x
8(x-1)=3x+7
2x+4(2x-1)=6(x+4)
i need help with these problem cant figure them out can someone please help me i really need thx to who ever can help me.

Answer by rfer(16322)

I will show you the first two.
x+5(x-1)=19
x+5x-5=19
6x=24
x=4
-----------------
2x=10(x-4)
2x=10x-40
-8x=-40
x=5